OpenStudy (elleblythe):

Find the solution set of the following inequality: (x+1)/(x+5) ≥[1/(x+5)]-x Please show complete solution and answer

OpenStudy (anonymous):

\[\dfrac{x+1}{x+5} \geq \dfrac{1}{x+5}-x\] You already have a common denominator - if you multiply each term by that denominator, you can eliminate them.

OpenStudy (elleblythe):

@gypsy1274 I'd like to see a final answer please so I could check if my own answer is right

OpenStudy (anonymous):

Sorry to be so suspicious, but so many on here are just looking for someone to do their homework for them. If you post your answer, I will post mine.

OpenStudy (elleblythe):

@gypsy1274 [-6, -5) U [0, +∞) ?

OpenStudy (anonymous):

I'm a bit rusty on the notation....Here is my answer: \(x\geq -6\)

OpenStudy (anonymous):

\(\left(\dfrac{x+1}{x+5} \geq \dfrac{1}{x+5}-x\right)\times (x+5)\) The denominators cancel with the first two terms leaving: \(x+1 \geq 1 - x^2-5x\)

OpenStudy (anonymous):

Then I moved everything over to the left side and factored.

OpenStudy (anonymous):

Is that what you did?

OpenStudy (anonymous):

\[x \ge 0\] \[-6 \le x < 5\]